OpenStudy (aryana_maria2323):

If 3x2 + y2 = 7 then evaluate the second derivative of y with respect to x when x = 1 and y = 2. Round your answer to 2 decimal places. Use the hyphen symbol, -, for negative values.

OpenStudy (anonymous):

start with \[6x+2yy'=0\] then solve for \(y'\) then repeat

OpenStudy (aryana_maria2323):

How would I do that? Can you please show me the steps? @satellite73

OpenStudy (anonymous):

algebra subtract \(6x\) from both sides, then divide both sides by \(2y\)

OpenStudy (aryana_maria2323):

so y= -6x/2y

OpenStudy (anonymous):

you might want to reduce

OpenStudy (aryana_maria2323):

wait is the answer. -6? I replaced x and y with the numbers given?

OpenStudy (anonymous):

no that is the first derivative, you need the second

OpenStudy (anonymous):

\[y'=-\frac{3x}{y}\] now you need \(y''\)

OpenStudy (anonymous):

this one requires the quotient rule

OpenStudy (aryana_maria2323):

Okay so what is y''?

OpenStudy (anonymous):

do you know the quotient rule?

OpenStudy (anonymous):

lets put the \(-3\) out front

OpenStudy (aryana_maria2323):

yes d/dx

OpenStudy (anonymous):

\[y'=-3\left(\frac{x}{y}\right)\]\[y''=-3\left(\frac{y-xy'}{y^2}\right)\]

OpenStudy (anonymous):

you good from there?

OpenStudy (aryana_maria2323):

Yes. I plugged in for x and y and I got -2.625 which rounds to -2.63 right?

OpenStudy (anonymous):

before you can plug in for x and y you have to replace \(y'\) by \(\frac{-3x}{y}\)

OpenStudy (aryana_maria2323):

yes I did and I got -1.5

OpenStudy (anonymous):

i believe you

OpenStudy (aryana_maria2323):

Haha so when you plug in y' and x and y I got -2.63. Is that correct?

OpenStudy (aryana_maria2323):

@satellite73

OpenStudy (aryana_maria2323):

You there @satellite73

OpenStudy (kayders1997):

Did you take the second derivative?

OpenStudy (aryana_maria2323):

I am no sure. I just did what ever satellite73 told me to do. is the second derivative y' ?

OpenStudy (kayders1997):

y''

OpenStudy (anonymous):

the first derivative is \(y'\) the second is \(y''\) but i get the feeling that it was not clear how to find them true or false?

OpenStudy (kayders1997):

after you find the first derivative you have to find the second one

OpenStudy (aryana_maria2323):

true. I found y' which was -1.5 and for y'' I got -2.625.

OpenStudy (anonymous):

you want to see how to find them?

OpenStudy (aryana_maria2323):

yes and is that not correct?

OpenStudy (anonymous):

i don't know, i didn't do the arithmetic but i can show you how to find the first and second derivative if you like

OpenStudy (aryana_maria2323):

Yes please. Once I find those is y'' my answer?

OpenStudy (anonymous):

looks good

OpenStudy (aryana_maria2323):

Yay okay so my answer is -2.63 when rounded?

OpenStudy (anonymous):

yes

OpenStudy (aryana_maria2323):

Thank you so much! Can you help me with one more question?

OpenStudy (anonymous):

sure if i can

OpenStudy (aryana_maria2323):

Find the slope of the graph of the relation x2y + 4y = 8 at the point (2, 1).

OpenStudy (noelgreco):

Differentiate implicitly.

OpenStudy (anonymous):

same thing only easier now lets see how to do it \(y\) is assumed to be a function of \(x\) so lets suppose it was \(y=\sin(x)\) what is the derivative of \[x^2\sin(x)+4\sin(x)\]?

OpenStudy (aryana_maria2323):

x^2cos(x)+2xsin(x)+4cos(x)

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

now lets replace \(\sin(x)\) by \(f(x)\) so \(\cos(x)=f'(x)\) what is the derivative of \[x^2f(x)+4f(x)\]?

OpenStudy (aryana_maria2323):

3x2f+4f

OpenStudy (anonymous):

hmm no think about this answer \[x^2\cos(x)+2x\sin(x)+4\cos(x) \] and how you used the product rule replace \)\sin(x)\) by \(f(x)\) and \(\cos(x)\) by \(f'(x)\)

OpenStudy (anonymous):

\[x^2\overbrace{\cos(x)}^{f'}+2x\overbrace{\sin(x)}^f+4\overbrace{\cos(x)}^{f'}\]

OpenStudy (aryana_maria2323):

Okay so x^2f'(x)+2xf(x)+4f'(x)

OpenStudy (anonymous):

ok now try i with \(y\) and \(y'\) instead of \(f(x)\) and \(f'(x)\)

OpenStudy (aryana_maria2323):

\[x^2y'+2xy+4y'\]

OpenStudy (anonymous):

right notice the use of the product rule

OpenStudy (anonymous):

and since the derivative of 8 is 0 you have \[x^2y'+2xy+4y'=0\] solve for \(y'\)

OpenStudy (aryana_maria2323):

I got y'=2x but I don't think that is right.

OpenStudy (aryana_maria2323):

you there @satellite73

OpenStudy (aryana_maria2323):

you there @satellite73

OpenStudy (anonymous):

oh hold on

OpenStudy (anonymous):

no, it is not right

OpenStudy (anonymous):

\[x^2y'+2xy+4y'=0\] solve for \(y'\) is like solving \[az+b+cz=0\] for \(z\)

OpenStudy (anonymous):

first subtract \(2xy\) get \[x^2y'+4y'=-2xy\] then combine like terms get \[(x^2+4)y'=-2xy\] then divide both sides by \(x^2+4\)

OpenStudy (aryana_maria2323):

That is where I messed up